the measures of two angles are in the ratio 4:5. the larger of the two angles is 60 less than twice the smaller angle. find the measure of each angle. complementary or supplementary?
Answer by cleomenius(959)

We will let x be the larger angle and y be the smaller angle.
angle y = 4/5x
Now, set up two equations
4x = 5y
x + 60 = 2y
4x - 5y = 0
x -2y = -60
-4x +8y = 240 Multiply the second equation by -4 and subtract from the first equation.
3y = 240
y = 80 degrees, which is 4/5 of x.
x = 100 degrees.
As a check, 2(80) - 60 = 100. This does check.
Cleomenius.